OpenStudy (richyw):

shortcut for solving a system of linear equations if I only need to solve for one of the variables?

OpenStudy (richyw):

so I need to solve this system of equations \[A-B-C=0\]\[C-D-F=0\]\[B+D-E=0\]\[2C+20F=1\]\[2C-3B+7D=0\]\[20F-7D-5E=0\] I can put these in an augmented matrix and row reduce to solve for all the variables, but the truth is I only need to find out what A is. What is the quickest way I can do this by hand?

OpenStudy (anonymous):

cramer's rule

OpenStudy (anonymous):

http://en.wikipedia.org/wiki/Cramer's_rule

OpenStudy (richyw):

I don't see how cramers rule will be any faster?

OpenStudy (anonymous):

all you have to do is solving two determinants .. as i see it it is much faster